# What If Moderation Didn’t Mean Suppression? A Case for Personalized Content Transformation

### v1

@@ 8. Limitations and Future Directions
Improving Selection Model. Our findings reveal that the effectiveness of an intervention is context-dependent, and that successful transformations manage a user’s cognitive state by providing closure. Our system’s selection framework, in contrast, relied on a limited user model derived from filter descriptions and lacked an explicit model of cognitive closure or the user’s immediate context. Improving its ability to select transformations that provide such closure is an important area for future work. The principles and preference data from this study provide a direct path to do so by aligning the selector with human preferences using alignment approaches such as DPO (Rafailov et al., 2023).
Long-Term Psychological Implications. Our study focused on users’ immediate experiences, which raises a longitudinal question: do personalized transformations support long-term wellbeing, or can they enable patterns of unhealthy avoidance? There is good reason to shield users from repeated triggers, as re-exposure may exacerbate anxiety (Markowitz and Fanselow, 2020). At the same time, indefinite avoidance may also have adverse psychological and social effects. Future work should explore how to design for this tension, perhaps by incorporating features that allow users to optionally and safely decrease intervention intensity over time, aligning the tool more closely with therapeutic goals.
Scope of Evaluation. Our user study, while providing qualitative insights, involved 15 participants primarily from Western contexts. Broader deployment is needed to understand potential cultural variations in how harm is perceived and what transformations are considered appropriate.
Technical and Platform Dependencies. While DIY-MOD’s core logic is generalizable, the content interception requires platform-specific adapters, and extending DIY-MOD to video-first (e.g., TikTok) platforms would require additional engineering effort. Further, our reliance on commercial VLMs means their capabilities, biases, and safety filters can affect performance, motivating future work on fine-tuning open-source models.
@@ A.1. Overview of Computational Load
The table’s first two groups, the Text Processing Sequence and the Image Filter Analysis, represent the critical path that must complete before returning an initial response to the client. The final Image Transformation Sequence can then process asynchronously in the background. Without our optimizations, this pipeline would require several minutes to complete; our approach reduces the critical path latency to 5-15 seconds.
@@ A.2. Asynchronous Processing with Multi-Batch Responses
Figure 4. DIY-MOD processing pipeline showing synchronous and asynchronous execution paths. The critical path (diagonal hatching) completes in 5-15 seconds, returning text modifications and polling tokens. Image transformations process asynchronously in background workers (dotted pattern). The client polls for completed transformations—missing on attempts until the content transformation is ready. Previously processed content can return immediately from cache on first poll.
@@ A.4. Content-Based Caching
$\text{cache\_key}=\text{hash}(\text{content})\oplus\text{hash}(\text{filter\_params})\oplus\text{hash}(\text{sensitivity\_level})$
@@ B.1. Data Analysis and Protection
We transcribed the interviews verbatim and then deleted the original recordings to avoid retaining identifiable data. Using reflexive thematic analysis (Braun and Clarke, 2019), we inductively identified the core themes presented in Section 3.2 through 3.7. We pseudonymized transcripts with coded identifiers (e.g., P_01, P_12). We stored names and email addresses for recruitment and compensation in a separate encrypted file, deleted them after payment, and never linked them to transcripts. All study data remained on encrypted, university-approved devices and cloud services accessible only to the core research team.
@@ C.1. DIYMOD
(4(a)) Chat interface for filter creation.
@@
(4(b)) Filter configuration interface. After establishing the filter description, users configure three parameters: (1) Content Type specifies whether the filter applies to text, images, or both; (2) Sensitivity Level indicates the user’s distress intensity; and (3) Duration sets filter expiration. For brevity, in popup interface we only showed three time presets. Users can modify the filter description anytime using the "Modify Filter" button or add metadata through the Options page for further customization.
@@ C.2. Dual Feed
Appendix D Intervention Description
Category Intervention Description & Relation to Design Space
Warning
Overlay
Stylistic Alteration*
Can apply to both sensitive region only or whole image
Intervention Description & Relation to Design Space
Overlay
Warning
Hides an entire text block behind a personalized warning message.

### v2

@@ 8. Limitations and Future Directions
Improving Selection Model. Our findings reveal that the effectiveness of an intervention is context-dependent, and that successful transformations manage a user’s cognitive state by providing closure. Our system’s selection framework, in contrast, relied on a limited user model derived from filter descriptions and lacked an explicit model of cognitive closure or the user’s immediate context. Improving its ability to select transformations that provide such closure is an important area for future work. The principles and preference data from this study provide a direct path to do so by aligning the selector with human preferences using alignment approaches such as DPO (rafailov2023direct).
Long-Term Psychological Implications. Our study focused on users’ immediate experiences, which raises a longitudinal question: do personalized transformations support long-term wellbeing, or can they enable patterns of unhealthy avoidance? There is good reason to shield users from repeated triggers, as re-exposure may exacerbate anxiety (markowitz2020exposure). At the same time, indefinite avoidance may also have adverse psychological and social effects. Future work should explore how to design for this tension perhaps by incorporating features that allow users to optionally and safely decrease intervention intensity over time, aligning the tool more closely with therapeutic goals.
Scope of Evaluation. Our user studies, while providing nuanced qualitative insights through a multi-stage design, involved a formative study with 12 participants, an in-situ evaluation with 15, and a preference elicitation with 12. This limited scale, along with a primarily Western participant pool, means broader deployment is needed to understand potential cultural variations in how harm is perceived and what transformations are considered appropriate.
Technical and Platform Dependencies. While DIY-MOD’s core logic is generalizable, the content interception requires platform-specific adapters, and extending DIY-MOD to video-first (e.g., TikTok) platforms would require additional engineering effort. Further, our reliance on commercial VLMs means their capabilities, biases, and safety filters can affect performance. We detail specific failure modes and mitigations in Section˜A.5; developing benchmarks to quantify edit failures and over-detection rates remains a potential future work.
Scalability & Latency. Our middleware approach also introduces computational and monetary costs. Our optimizations (detailed in Appendix˜A) keep the per-batch latency within a usable range for real-time interaction, but high cost associated with LLM/VLM usage remains a scalability constraint. These constraints highlight the efficiency gains of a platform-integrated approach (Section˜7.2). Future work could also explore smaller, specialized on-device models to reduce both latency and cost.
@@
Misuse Potential. The same open-ended agency that enables protective filtering could be inverted to highlight content that might cause harm to users themselves. This risk is inherent to any system that provides users with fine-grained control over content presentation. In practice, such misuse is bounded by the safety constraints of the underlying VLM. The system cannot execute interventions that the foundation model itself refuses to generate.
@@ A.1. Overview of Computational Load
The table’s first two groups, the Text Processing Sequence and the Image Filter Analysis, represent the critical path that must complete before returning an initial response to the client for the full batch of posts. The final Image Transformation Sequence can then process asynchronously in the background generating time-consuming image transformation and scoring. Without any optimizations, this pipeline would require several minutes to complete; our approach reduces the critical path latency to 5-15 seconds for the batch.
@@ A.2. Asynchronous Processing with Multi-Batch Responses
Figure 5. DIY-MOD processing pipeline showing synchronous and asynchronous execution paths. The critical path (diagonal hatching) completes in 5-15 seconds for a batch of posts, returning text modifications and polling tokens. Image transformations process asynchronously in background workers (dotted pattern). The client polls for completed transformations—missing on attempts until the content transformation is ready. Previously processed content can return immediately from cache on first poll. Note that, the critical path delay is for a batch of usually 25 posts in reddit. And, this delay only happens when the user first loads/refreshes the feed in browser. For subsequent batches of posts within the browsing session, predictive prefetching (Section˜A.3) minimizes the delay under reasonable scrolling behavior.
@@ A.4. Content-Based Caching
$\mathtt{cache\_key}=\mathtt{hash}(\mathtt{content})\oplus\mathtt{hash}(\mathtt{filter\_params})\oplus\mathtt{hash}(\mathtt{sensitivity\_level})$
@@
A.5. Edit Failures, False Obfuscations and Mitigation Strategies
A.5.1. Generation failures and safeguards.
During deployment, we encountered two types of failures in our generation pipeline. First, specific filters occasionally triggered broader VLM categorization: for example, a filter for “Middle East Conflict” by one participant matched general conflict imagery or rubble. Second, Gemini’s safety guardrails sometimes refused to generate lawful sensitive content such as medical imagery. We investigated this further but observed non-reproducible failures where identical prompts and inputs would succeed on some attempts but fail on others.
We designed our mitigation strategy to treat both as fail-safe defaults. When generation fails, we first retry, which resolve in most cases. Generating $K=3$ parallel candidates also increases the likelihood that at least one succeeds and scored higher. While we could switch to alternative VLMs upon multiple failure, we prioritized real-time feed rendering over exhaustive fallback strategies. Even when all candidates fail to generate, we leave the content unchanged but still apply the transparency indicator (Section˜4.2, Figure˜1(d)) on top of the original post. This indicator informs users that the content matched one of their filters and was meant to be transformed. Participants reported using this indicator as a personalized warning label—a cue to look away from posts that might be harmful to them. In other words, we err on the side of flagging a post as potentially triggering rather than exposing users to it.
A.5.2. Audit of edit failures and false positives.
To quantify edit failures and false obfuscations, we manually audited posts that our extension marked as modified via the transparency indicator. For each participant, we randomly sampled 5 text posts and 5 image posts, yielding 75 text items and 75 image items. For text, we confirmed via screen recordings that all 75 posts were successfully transformed. For images, we visually matched each transformed image to the original Reddit post and the filter description of the viewer. We found that 66/75 ($\approx 88\%$) images were visually transformed, while 8 ($\approx 12\%$) were either left unchanged or edited in the wrong region. Among the 66 transformed images, 62 ($\approx 93\%$) were correctly transformed according to the viewer’s filter description, implying a false-positive transformation rate of roughly 7% of transformed images.
@@ B.1. Data Analysis and Protection
We transcribed the interviews verbatim and then deleted the original recordings to avoid retaining identifiable data. Using reflexive thematic analysis (braun2019reflecting), we inductively identified the core themes presented in Section 3.2 through 3.7. We pseudonymized transcripts with coded identifiers (e.g., P_01, P_12). We stored names and email addresses for recruitment and compensation in a separate encrypted file, deleted them after payment, and never linked them to transcripts. All study data remained on encrypted, university-approved devices and cloud services accessible only to the core research team.
@@ C.1. DIYMOD
(5(a)) Chat interface for filter creation.
@@
(5(b)) Filter configuration interface. After establishing the filter description, users configure three parameters: (1) Content Type specifies whether the filter applies to text, images, or both; (2) Sensitivity Level indicates the user’s distress intensity; and (3) Duration sets filter expiration. For brevity, in popup interface we only showed three time presets. Users can modify the filter description anytime using the "Modify Filter" button or add metadata through the Options page for further customization.